Given that the ionization constant for water (Kw) is 4.52x10-15 at 15 °C, calculate the pH for pure water at 15 °C Report your answer to 3 decimal places. 16.5

Not sure if this is right, pure water so its [H] squared

Kw = [H] squared so I square rooted and took the -log of this to get the pH
Answer #1

Kw = [H+][OH-]

since water is neutral,

[H+] = [OH-]

So,

Kw = [H+][H+]

4.52*10^-15 = [H+]^2

[H+] = 6.72*10^-8 M

use:

pH = -log [H+]

= -log (6.72*10^-8)

= 7.1726

Answer: 7.173